OpenStudy (anonymous):

Given the function f(x) = log3(x + 4), find the value of f−1(3).

OpenStudy (solomonzelman):

You want to find, \(\large\color{black}{ \displaystyle f^{-1}\left(3\right) }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y=\log_3(x+4)}\)

OpenStudy (solomonzelman):

\(\bf [1]\) swipe around x and y. \(\bf [2]\) isolate the y. \(\bf [{~}lastly{~}]\) plug in 3 for x.

OpenStudy (solomonzelman):

Or, you can simply plug in 3 for y, and solve for x.

OpenStudy (solomonzelman):

\(\large\color{red}{3}=\log_3(x+4)\)

OpenStudy (anonymous):

so I swipe x and y in the first equation?

OpenStudy (solomonzelman):

Yes, you can swipe x and y in the first equation then solve for x, and then plug in 3 for x.

OpenStudy (solomonzelman):

Or, alternatively, you can plug in 3 for y, and solve for x.

OpenStudy (solomonzelman):

Either way you get the answer, except that way 1 is standard and way 2 is quicker.

OpenStudy (anonymous):

x=23 correct?

OpenStudy (solomonzelman):

I don't really know.... can you please tell me, how did you get this result?

OpenStudy (anonymous):

substituted 3 for y

OpenStudy (solomonzelman):

yes, and then solve for x. 23 is correct

OpenStudy (solomonzelman):

So you will get that: \(\large\color{black}{ \displaystyle f^{-1}(3)=23 }\)

OpenStudy (anonymous):

Thank you for your help!!

OpenStudy (solomonzelman):

yw